Shirley is currently five times as old as Harry. In six years, Shirley will be twice as old as Harry.
a.How old are they now?
b. How old will they be in 6 years?

Answer by stanbon(75887)

Shirley is currently five times as old as Harry. In six years, Shirley will be twice as old as Harry.
---
Let Harry be "h" NOW
Then Shirley is "5h" NOW.
-------------------------------
Equation:
(5h+6) = 2(h+6)
5h + 6 = 2h + 12
3h = 6
h = 2 (Harry's age NOW)
5h = 10 (Shirley's age NOW)
=================================
a.How old are they now?
=================================
b. How old will they be in 6 years?
h+6 = 8 (Harry's age in 6 years)
5h+6 = 16 (Shirley's age in 6 years)
==================================
Cheers,
Stan H.